### v1

@@ 1 Introduction
Reconfigurable Intelligent Surfaces (RISs) have become a focal point in ISAC research due to their ability to dynamically shape the propagation environment and extend the coverage of the system [12]. Integrating RISs into ISAC systems offers a new dimension in system design, as RIS placement influences available sensing points, thus improving coverage and resolution [13, 2]. However, balancing the performance tradeoffs between sensing and communication in RIS-aided ISAC systems is a critical concern [14]. Existing approaches often treat the RIS as an extension of the base station (BS), focusing on the joint beampattern design for both user coverage and sensing [15, 16, 17]; these joint designs can become increasingly complex, leading to sequential design strategies in which BS and RIS configurations are optimized separately [16, 17]. Another challenge is represented by the heavy attenuation of the indirect link generated by the RIS: this may compromise the sensing function in challenging situations, such as those of detecting dim moving targets like small drones, or the communication QoS, if an increased fraction of power is dedicated to the sensing function.
@@
In this paper, we consider an OFDM-based ISAC system where a BS serves multiple ground users while using a nearby RIS for detecting non-cooperative airborne targets. The main contribution is the integration of the multi-frame TBD processor from [22, 23, 24] into the design of the ISAC transceiver in order to exploit the temporal correlation of the echoes generated by a moving target. Furthermore, we define the two-way beampattern of the considered RIS-assisted radar, which specifically accounts for the underlying OFDM-based signal structure and the effects of both the transmit and receive beampatterns. A design method is proposed for the two-way beampattern, that maximizes the expected power for any target echo within the inspected subvolume, while controlling the transmit and/or receive sidelobe levels in the presence of eavesdroppers and/or interference sources and avoiding any impact on the communication function. Finally, we discuss some numerical examples based on realistic 5G system specifications. Remarkably, our results demonstrate that increasing the number of scans processed by the radar detector (i.e., increasing its implementation complexity) allows for reduced power dedicated to the radar function, while maintaining the same sensing performance (measured in terms of probability of target detection and root mean square error in the estimation of the target position); this excess power can then be redirected to the communication function, so as to increase the user sum-rate.
Previous related studies have already investigated several strategies for designing the downlink spatial beamformers of the BS [26, 27, 28, 29, 30, 31]. In particular, relevant performance metrics for the communication function are the signal-to-interference-plus-noise-ratio (SINR) and the rate of the users, while the signal-to-noise-plus-clutter ratio, the Cramér Rao lower bound on the estimation of the unknown target parameters, and the transmit beampattern are common metrics for the sensing function. Common design criteria aim to maximize a communication metric under a constraint on a sensing metric [26, 27, 28, 29] or vice-versa [30, 31]. This contribution significantly differs from previous studies in implementation of the sensing function, as outline is Table 1. For the first time, we incorporate multi-frame radar processing into a RIS-aided ISAC transceiver and recognize it as a means to obtain more favorable sensing and communication tradeoffs. Also, for given downlink communication beamformers (that can be obtained for example from the aforementioned existing methods), we show that the RIS response and the sensing beamformer used by the BS to illuminate the RIS can be jointly designed to scan the volume of interest while controlling the sidelobes of the transmit and receive beampatterns and avoiding any interference on the communication function.
@@
Organization: The remainder of this work is organized as follows. Sec. 2 presents the system description and the underlying design assumptions. Sec. 3 illustrates the proposed system design. Sec. 4 contains some numerical examples. Finally, concluding remarks are given in Sec. 5.
Notation: Column vectors and matrices are denoted by lowercase and uppercase boldface letters, respectively. The symbols $(\cdot)^{\mathsf{T}}$, $(\cdot)^{*}$, $(\cdot)^{\mathsf{H}}$ and $(\cdot)^{+}$ denote the transpose, conjugate, conjugate-transpose and pseudo-inverse operations, respectively. $\bm{I}_{M}$ is the $M\times M$ identity matrix, while $\bm{0}_{M}$ is an $M\times 1$ vector of all zero elements. $\|\bm{\alpha}\|$ is the Euclidean norm of the vector $\bm{\alpha}$ and $\text{diag}\{\bm{\alpha}\}$ is the $N\times N$ diagonal matrix with the entries of vector $\bm{\alpha}$ on the main diagonal. $[\bm{A}]_{i,j}$ marks the element in the $i$-th row and $j$-th column for matrix $\bm{A}$. Finally, the symbols $\mathrm{i}$, $\varnothing$ and $\mathbb{E}\{\cdot\}$ denote the imaginary unit, an empty set, and the statistical expectation, respectively.
@@ 2 System description
We assume that the users are not in the field-of-view of the RIS due to the specific system geometry and/or the presence of a blockage. Also, we make a narrowband assumption on each subcarrier, whereby $(\Delta_{\rm tx}+\Delta_{\rm rx}+\Delta_{\rm ris})\ll c_{o}/W_{o}$, where $\Delta_{\rm tx}$, $\Delta_{\rm rx}$, and $\Delta_{\rm ris}$ are the largest distance among any element pair in the transmit array, receive array, and RIS, respectively, and $c_{o}$ is the speed of light [33]. Finally, let $\tau_{o}$ be the maximum delay over any BS-user or BS-RIS-target-RIS-BS path; we assume $\tau_{o}\leq T_{o}$, so that no interference among subsequent OFDM symbols is present at the user and radar receivers.
The volume $\mathcal{V}_{r}$ is in the RIS far-field and, when observed from a Cartesian reference system at the center of gravity of the RIS, spans the range, azimuth, and elevation intervals $[R_{\min},R_{\max}]$, $[\theta_{\min}^{\rm{az}},\theta_{\max}^{\rm{az}}]$, and $[\theta_{\min}^{\rm{el}},\theta_{\max}^{\rm{el}}]$, respectively; [^1] The RIS is located on the $(x,z)$-plane with its elements reflecting towards the positive $y$-axis. The azimuth angle of a point $\bm{p}\in\mathbb{R}^{3}$ is the angle between the $y$-axis and the orthogonal projection of the vector pointing towards $\bm{p}$ onto the $(x,y)$-plane, which is positive when going anti-clockwise from the $y$-axis; also, the elevation angle is the angle between the vector pointing towards $\bm{p}$ and its orthogonal projection onto the $(x,y)$-plane, which is positive when going towards the positive $z$-axis from the $(x,y)$-plane. the far-field assumption implies that $R_{\min}\geq 2\Delta_{\rm ris}^{2}/\lambda_{-}$, where $\lambda_{-}=c_{o}/(f_{o}-0.5N_{o}W_{o})$ is the minimum wavelength [34]. Let $\bm{\theta}=[\theta^{\rm az}\;\theta^{\rm el}]^{\mathsf{T}}$ be the angular direction specified by azimuth and elevation angles $\theta^{\rm az}$ and $\theta^{\rm el}$, respectively; then, we denote by $\bm{t}_{q}(\bm{\theta})$ the far-field steering vector of the RIS towards $\bm{\theta}$ on the $q$-th subcarrier, for $q=1,\ldots,N_{\rm sub}$.
Figure 2: Scanning policy for the considered ISAC scenario.
@@ 2.2 Radar received signal
We assume that the channel matrices $\bm{G}_{q,\rm rx}$ and $\bm{G}_{q,\rm tx}$ remain constant over the CPI and are known to the BS; for example, they can be estimated in the radar idle periods. In the considered CPI, we define the two-way (power) beampattern of the RIS-aided radar on the $q$-th subcarrier as
@@
where $\text{BP}_{q,{\rm tx}}(\bm{\theta})$ and $\text{BP}_{q,{\rm rx}}(\bm{\theta})$ are the transmit and receive beampatterns, respectively. We also define the two-way, transmit, and receive beampatterns of the RIS-aided radar over the entire frequency band as
@@ 2.3 User received signal
for $k=1,\ldots,K$, $q=1,\ldots,N_{\rm sub}$, and $n=1,\ldots,N_{\rm sym}$, where $\bm{h}_{q,k}\in\mathbb{C}^{D_{\rm tx}}$ is the channel vector between the BS and the $k$-th user on the $q$-th subcarrier and $z_{q,k}(n)\in\mathbb{C}$ is the additive noise, modeled as a circularly-symmetric Gaussian random variable with variance $\sigma^{2}_{k}$.
Let $\bm{H}_{q}=[\bm{h}_{q,1}\,\cdots\,\bm{h}_{q,K}]\in\mathbb{C}^{D_{\rm tx}\times K}$ be the downlink channel matrix in the $q$-th subcarrier. We assume that $\bm{H}_{q}$ is full rank, constant over the CPI, and known to the BS: for example, this matrix can be estimated in the radar idle periods. Accordingly, the SINR of the $k$-th user on the $q$-th subcarrier is
@@
for $k=1,\ldots,K$ and $q=1,\ldots,N_{\rm sub}$. Assuming Gaussian signaling, the expression $\log_{2}(1+\mathrm{SINR}_{q,k})$ indicates the number of bits that can be reliably transmitted to user $k$ on the $q$-th subcarrier within one OFDM symbol. By summing these expressions over the subcarrier index $q$, we can thus determine the achievable rate (in bits per OFDM symbol) for user $k$ as follows:
@@
for $k=1,\ldots,K$. The corresponding network sum-rate is
$\mathcal{R}=\sum_{k=1}^{K}\mathcal{R}_{k},$ (10)
still measured in terms of reliably transmitted bits per OFDM symbol. Notice that the rate of each user depends on the BS power vectors $\{\bm{\gamma}_{q}\}_{q=1}^{N_{\rm{sub}}}$ and the BS precoding matrices $\{\bm{F}_{q}\}_{q=1}^{N_{\rm{sub}}}$ (more on this in Sec. 3).
@@ 3 System design
    for $q=1,\ldots,N_{\rm sub}$. Also, the fraction of power allocated to each user on the $q$-th subcarrier and the corresponding beamformers, namely, $\{\gamma_{q,k}\}_{k=1}^{K}$ and $\{\bm{f}_{q,k}\}_{k=1}^{K}$, are chosen to optimize the communication function, for $q=1,\ldots,N_{\rm sub}$: any of the design criteria already devised in the literature can be employed here: [^2] Notice in passing that meaningful designs usually ensure that the vectors $\{\bm{f}_{q,k}\}_{k=1}^{K}$ belong to the column-space of $\bm{H}_{q}$, as any power sent outside such linear subspace will not be received by the users. for example, see [35, 36].
@@ 3.1 Design of the radar beamformers and RIS response
$\max_{\begin{subarray}{c}\bm{\omega}\in\Omega^{D_{\rm ris}},\\ \{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}\in\mathbb{C}^{D_{\rm tx}-K}\end{subarray}}$ $\sum_{\bm{\theta}\in\Theta_{i}}\text{BP}\left(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}\right),$
s.t. $\;\text{BP}\left(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}\right)\leq\varepsilon_{\rm sl},\forall\bm{\theta}\in\Theta_{i,{\rm sl}},$
$\;\text{BP}_{\rm{tx}}\left(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}\right)\leq\varepsilon_{\rm ev},\forall\bm{\theta}\in\Theta_{\rm ev},$
@@
where the dependence on the optimization variables $\bm{\omega}$ and $\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}$ has been made explicit in the two-way, transmit, and receive beampatterns; in (3.1), $\varepsilon_{\rm sl},\varepsilon_{\rm ev},\varepsilon_{\rm ja}$ are upper bounds to the two-way beampattern in the sidelobe region, to the transmit beampattern in the eavesdropper region, and to the receive beampattern to the jammer region, while the last constraint forces the radar beamformers to be unit-modulus.
@@
Problem (3.1) is non-convex and mixed-integer. Upon resorting to the penalty method, the constrained optimization problem is replaced by a series of one or more unconstrained problems obtained by penalizing the objective function with an extra cost (penalty function) that discourages the violations of the constraints. In particular, we approximate (3.1) as [^3] Notice that we can get arbitrary close to the optimal objective value of (3.1) by solving (3.1) with $\beta$ sufficiently large [40, Th. 9.2.2].
@@
$\max_{\begin{subarray}{c}\bm{\omega}\in\Omega^{D_{\rm ris}},\\ \{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}\in\mathbb{C}^{D_{\rm tx}-K}\end{subarray}}$$\Biggl{\{}\sum_{\bm{\theta}\in\Theta_{i}}\text{BP}\left(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}\right)$
@@
-- $\beta\Biggl{[}\sum_{\bm{\theta}\in\Theta_{i,{\rm sl}}}g\left(\text{BP}\left(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}\right)-\varepsilon_{\rm sl}\right)$
$+$ $\sum_{\bm{\theta}\in\Theta_{\rm ev}}g\left(\text{BP}_{\rm{tx}}\left(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}\right)-\varepsilon_{\rm ev}\right)$
$+$ $\sum_{\bm{\theta}\in\Theta_{\rm ja}}g\left(\text{BP}_{\rm{rx}}\left(\bm{\theta};\bm{\omega}\right)-\varepsilon_{\rm ja}\right)\Biggr{]}\Biggr{\}},$
s.t. $\;\|\bm{U}_{q}\tilde{\bm{f}}_{q,r}\|^{2}=1,\;q=1,\ldots,N_{\rm sub},$ (13)
where $\beta>0$ is a conveniently large number, and $g(\chi)=(\max\{0,\chi\})^{2}$ is the standard quadratic penalty function [40].
@@
Problem (3.1) is still quite complex, and we propose to find a suboptimal solution by resorting to the block-coordinate ascent method [40]: starting from a feasible point, the objective function is maximized with respect to each of the block of variables, taken in cyclic order, while keeping the other ones fixed at their previous values. The block variables are $\omega_{1}$, $\omega_{2}$,…, $\omega_{D_{\rm ris}}$, and $\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}$. Therefore, we are faced with $D_{\rm ris}+1$ reduced-complexity sub-problems: $D_{\rm ris}$ sub-problems for the maximization over each RIS phase and one sub-problem for the maximization over the radar beamformers.
@@
As to the RIS phase $\omega_{d}$, a suboptimal solution can be found by resorting to the projected gradient algorithm. In particular, the gradient of the objective function of Problem (3.1) is
$\bm{\nabla}_{\rm ris}$ $=2\sum_{\bm{\theta}\in\Theta_{i}}\sum_{q=1}^{N_{\rm sub}}\Bigl{(}\bm{\omega}^{\mathsf{H}}\bm{B}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}\bm{A}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}$
$\quad+\bm{\omega}^{\mathsf{H}}\bm{A}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}\bm{B}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}\Bigr{)}$
@@
$\quad-2\beta\Bigg{[}\sum_{\bm{\theta}\in\Theta_{i,{\rm sl}}}g^{\prime}\left(\text{BP}\left(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}\right)-\varepsilon_{\rm sl}\right)$
×$\quad\times\sum_{q=1}^{N_{\rm sub}}\Bigl{(}\bm{\omega}^{\mathsf{H}}\bm{B}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}\bm{A}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}$
$\quad+\bm{\omega}^{\mathsf{H}}\bm{A}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}\bm{B}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}\Bigr{)}$

### v2

@@ 1 Introduction
Reconfigurable Intelligent Surfaces (RISs) have become a focal point in ISAC research due to their ability to dynamically shape the propagation environment and extend the coverage of the system [12]. Integrating RISs into ISAC systems offers a new dimension in system design, as RIS placement influences available sensing points, thus improving coverage and resolution [13, 2]. However, balancing the performance tradeoffs between sensing and communication in RIS-aided ISAC systems is a critical concern [14]. Existing approaches often treat the RIS as an extension of the base station (BS), focusing on the joint beampattern design for both user coverage and sensing [15, 16, 17, 18]; these joint designs can become increasingly complex, leading to sequential design strategies in which BS and RIS configurations are optimized separately [16, 17, 18]. A key challenge in RIS-aided systems is obtaining accurate channel state information (CSI) for all RIS-to-user channels, which requires more elaborated training procedures [19, 20, 21]. Another challenge is the multiplicative path loss in the indirect link generated by the RIS, which may be mitigated by deploying the RIS closer to the BS and/or by resorting to active elements that amplify the reflected signal [22].
@@
1.1 Motivations and Contributions
Driven by the growing demand to integrate sensing and communication capabilities, this paper explores an OFDM-based ISAC system where a BS serves multiple ground users, while a nearby RIS is exploited for detecting non-cooperative airborne targets. The main contributions of our work are summarized below.
  • We integrate the multi-frame TBD processor from [27, 28, 29] into the ISAC transceiver design in order to exploit the temporal correlation of the echoes generated by a moving target.
  • We define the two-way beampattern of the considered RIS-assisted radar, which specifically accounts for the underlying OFDM-based signal structure and the effects of both the transmit and receive beampatterns.
  • We propose a method to design the two-way beampattern, that maximizes the expected power for any target echo within the inspected subvolume, while controlling the transmit and/or receive sidelobe levels in the presence of eavesdroppers and/or interference sources and avoiding any impact on the communication function.
  • Finally, we provide numerical results based on realistic 5G system specifications. Remarkably, our results demonstrate that increasing the number of scans processed by the radar detector (i.e., increasing its implementation complexity) allows for reduced power dedicated to the radar function, while maintaining the same sensing performance (measured in terms of probability of target detection and root mean square error in the estimation of the target position); this excess power can then be redirected to the communication function, so as to increase the user sum-rate.
Previous related studies have already investigated several strategies for designing the downlink spatial beamformers of the BS [31, 32, 33, 34, 35, 36]. In particular, relevant performance metrics for the communication function are the signal-to-interference-plus-noise-ratio (SINR) and the rate of the users, while the signal-to-noise-plus-clutter ratio, the Cramér Rao lower bound on the estimation of the unknown target parameters, and the transmit beampattern are common metrics for the sensing function. Common design criteria aim to maximize a communication metric under a constraint on a sensing metric [31, 32, 33, 34] or vice-versa [35, 36]. This paper significantly differs from previous studies in the implementation of the sensing function, as outlined is Table 1. To the best of our knowledge, this is the first work to incorporate a multi-frame radar processing into an ISAC transceiver, recognizing it as a means to obtain more favorable sensing and communication tradeoffs. Also, for given downlink communication beamformers (that can be obtained for example from the aforementioned existing methods), we show that the RIS response and the sensing beamformer used by the BS to illuminate the RIS can be jointly designed to scan the volume of interest while controlling the sidelobes of the transmit and receive beampatterns and avoiding any interference on the communication function.
@@ 1.1 Motivations and Contributions
[16] $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$
[17] $\times$ $\times$ $\times$ $\checkmark$ $\times$ $\times$ $\checkmark$
@@
1.2 Organization
The remainder of this work is organized as follows. Sec. 2 presents the system description and the underlying design assumptions. Sec. 3 illustrates the proposed beampattern design and the incorporation of the TBD processor into the RIS-assisted radar transceiver. Sec. 4 contains some numerical examples exploring the achievable tradeoffs between the communication and sensing performance. Finally, concluding remarks are given in Sec. 5.
1.3 Notation
Column vectors and matrices are denoted by lowercase and uppercase boldface letters, respectively. The symbols $(\cdot)^{\mathsf{T}}$, $(\cdot)^{*}$, $(\cdot)^{\mathsf{H}}$ and $(\cdot)^{+}$ denote the transpose, conjugate, conjugate-transpose and pseudo-inverse operations, respectively. $\bm{I}_{M}$ is the $M\times M$ identity matrix, while $\bm{0}_{M}$ is an $M\times 1$ vector of all zero elements. $\|\bm{\alpha}\|$ is the Euclidean norm of the vector $\bm{\alpha}$ and $\text{diag}\{\bm{\alpha}\}$ is the $N\times N$ diagonal matrix with the entries of vector $\bm{\alpha}$ on the main diagonal. $[\bm{A}]_{i,j}$ marks the element in the $i$-th row and $j$-th column for matrix $\bm{A}$. Finally, the symbols $\mathrm{i}$, $\varnothing$ and $\mathbb{E}\{\cdot\}$ denote the imaginary unit, an empty set, and the statistical expectation, respectively.
@@ 2 System description
We assume that the users are not in the field-of-view of the RIS due to the specific system geometry and/or the presence of a blockage. [^1] The following developments can be adapted to a different scenario where the RIS also supports the communication function. For example, if the RIS is partitioned into two sub-arrays [38], one sub-array can be dedicated to the sensing function and the other one to the communication function. Also, we make a narrowband assumption on each subcarrier, whereby $(\Delta_{\rm tx}+\Delta_{\rm rx}+\Delta_{\rm ris})\ll c_{o}/W_{o}$, where $\Delta_{\rm tx}$, $\Delta_{\rm rx}$, and $\Delta_{\rm ris}$ are the largest distance among any element pair in the transmit array, receive array, and RIS, respectively, and $c_{o}$ is the speed of light [39]. Finally, let $\tau_{o}$ be the maximum delay over any BS-user or BS-RIS-target-RIS-BS path; we assume $\tau_{o}\leq T_{o}$, so that no interference among subsequent OFDM symbols is present at the user and radar receivers.
The volume $\mathcal{V}_{r}$ is in the RIS far-field and, when observed from a Cartesian reference system at the center of gravity of the RIS, spans the range, azimuth, and elevation intervals $[R_{\min},R_{\max}]$, $[\theta_{\min}^{\rm{az}},\theta_{\max}^{\rm{az}}]$, and $[\theta_{\min}^{\rm{el}},\theta_{\max}^{\rm{el}}]$, respectively; the far-field assumption implies that $R_{\min}\geq 2\Delta_{\rm ris}^{2}/\lambda_{-}$, where $\lambda_{-}=c_{o}/(f_{o}-0.5N_{o}W_{o})$ is the minimum wavelength [40]. Let $\bm{\theta}=[\theta^{\rm az}\;\theta^{\rm el}]^{\mathsf{T}}$ be the angular direction specified by azimuth and elevation angles $\theta^{\rm az}$ and $\theta^{\rm el}$, respectively; then, we denote by $\bm{t}_{q}(\bm{\theta})$ the far-field steering vector of the RIS towards $\bm{\theta}$ on the $q$-th subcarrier, for $q=1,\ldots,N_{\rm sub}$.
@@ 2.1 Transmit signal
Figure 2: Scanning policy for the considered ISAC scenario.
@@ 2.2 Radar received signal
We assume that the channel matrices $\bm{G}_{q,\rm rx}$ and $\bm{G}_{q,\rm tx}$ remain constant over the CPI and are known to the BS; for example, they can be estimated in the radar idle periods. [^2] The BS-to-RIS channel changes over long time scales when both BS and RIS are mounted on top of tall structures: this implies that an accurate estimate of the BS-to-RIS channel can be obtained in practice. In the considered CPI, we define the two-way (power) beampattern of the RIS-aided radar on the $q$-th subcarrier as
@@
where $\text{BP}_{q,{\rm tx}}(\bm{\theta})$ and $\text{BP}_{q,{\rm rx}}(\bm{\theta})$ are the transmit and receive beampatterns, which accounts for the power transmitted towards $\bm{\theta}$ and the power received from $\bm{\theta}$, respectively. We also define the two-way, transmit, and receive beampatterns of the RIS-aided radar over the entire frequency band as
@@ 2.3 User received signal
for $k=1,\ldots,K$, $q=1,\ldots,N_{\rm sub}$, and $n=1,\ldots,N_{\rm sym}$, where $\bm{h}_{q,k}\in\mathbb{C}^{D_{\rm tx}}$ is the channel vector between the BS and the $k$-th user on the $q$-th subcarrier and $z_{q,k}(n)\in\mathbb{C}$ is the additive noise, modeled as a circularly-symmetric Gaussian random variable with variance $\sigma^{2}_{k}$. In the $q$-th subcarrier, let $\bm{H}_{q}=[\bm{h}_{q,1}\,\cdots\,\bm{h}_{q,K}]\in\mathbb{C}^{D_{\rm tx}\times K}$ be the downlink channel matrix. We assume that $\bm{H}_{q}$ is full rank, constant over the CPI, and known to the BS: for example, this matrix can be estimated in the radar idle periods. [^3] To simplify exposition, perfect CSI between BS and users is assumed in this study. Indeed, designing the communication beamformers at the BS and assessing the impact of possible channel estimation errors on the communication performance is not main focus of this study. Accordingly, the SINR of the $k$-th user on the $q$-th subcarrier is
@@
for $k=1,\ldots,K$ and $q=1,\ldots,N_{\rm sub}$. Assuming Gaussian signaling, the number of bits that can be reliably transmitted to user $k$ on the $q$-th subcarrier within one OFDM symbol is $\log_{2}(1+\mathrm{SINR}_{q,k})$; hence, the achievable rate (in bits per OFDM symbol) for user $k$ is
@@
for $k=1,\ldots,K$. Finally, the user sum-rate is
$\mathcal{R}=\sum_{k=1}^{K}\mathcal{R}_{k}.$ (10)
Notice that the rate of each user depends on the BS power vectors $\{\bm{\gamma}_{q}\}_{q=1}^{N_{\rm{sub}}}$ and the BS precoding matrices $\{\bm{F}_{q}\}_{q=1}^{N_{\rm{sub}}}$ (more on this in Sec. 3).
@@ 3 System design
    for $q=1,\ldots,N_{\rm sub}$. Also, the fraction of power allocated to each user on the $q$-th subcarrier and the corresponding beamformers, namely, $\{\gamma_{q,k}\}_{k=1}^{K}$ and $\{\bm{f}_{q,k}\}_{k=1}^{K}$, are chosen to optimize the communication function, for $q=1,\ldots,N_{\rm sub}$: any of the design criteria already devised in the literature can be employed here: [^4] Notice in passing that meaningful designs usually ensure that the vectors $\{\bm{f}_{q,k}\}_{k=1}^{K}$ belong to the column-space of $\bm{H}_{q}$, as any power sent outside such linear subspace will not be received by the users. for example, see [41, 42, 43].
@@ 3.1 Design of the radar beamformers and RIS response
$\max_{\begin{subarray}{c}\bm{\omega}\in\Omega^{D_{\rm ris}},\\ \{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}\\ \in\mathbb{C}^{D_{\rm tx}-K}\end{subarray}}$ $\sum_{\bm{\theta}\in\Theta_{i}}\text{BP}\left(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}\right),$ (12a)
s.t. $\;\text{BP}\left(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}\right)\leq\varepsilon_{\rm sl},\forall\bm{\theta}\in\Theta_{i,{\rm sl}},$ (12b)
$\;\text{BP}_{\rm{tx}}\!\left(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}\right)\!\leq\!\varepsilon_{\rm ev},\forall\bm{\theta}\in\Theta_{\rm ev},$ (12c)
@@
where the dependence on the optimization variables $\bm{\omega}$ and $\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}$ has been made explicit in the two-way, transmit, and receive beampatterns; in Problem (12), $\varepsilon_{\rm sl},\varepsilon_{\rm ev},\varepsilon_{\rm ja}$ are upper bounds to the two-way beampattern in the sidelobe region, to the transmit beampattern in the eavesdropper region, and to the receive beampattern to the jammer region, while the last constraint forces the radar beamformers to be unit-modulus.
@@
Problem (12) is non-convex and mixed-integer. Upon resorting to the penalty method, the constrained optimization problem is replaced by a series of one or more unconstrained problems obtained by penalizing the objective function with an extra cost (penalty function) that discourages the violations of the constraints. In particular, we propose to approximate Problem (12) with Problem (13) reported at the top of the next page, where $\beta>0$ is a conveniently large number and $g(\chi)=(\max\{0,\chi\})^{2}$ is the standard quadratic penalty function [47]. [^5] Notice that we can get arbitrary close to the optimal objective value of (12) by solving (13) with $\beta$ sufficiently large [47, Th. 9.2.2].
@@
$\max_{\begin{subarray}{c}\bm{\omega}\in\Omega^{D_{\rm ris}},\\ \{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}\in\mathbb{C}^{D_{\rm tx}-K}\end{subarray}}$$\Biggl{\{}\sum_{\bm{\theta}\in\Theta_{i}}\text{BP}(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}})-\beta\Biggl{[}\sum_{\bm{\theta}\in\Theta_{i,{\rm sl}}}g\Bigl{(}\text{BP}(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}})-\varepsilon_{\rm sl}\Bigr{)}\Biggr{]}$
@@
$-\beta\Biggl{[}\sum_{\bm{\theta}\in\Theta_{\rm ev}}g\Bigl{(}\text{BP}_{\rm{tx}}(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}})-\varepsilon_{\rm ev}\Bigr{)}\Biggr{]}-\beta\Biggl{[}\sum_{\bm{\theta}\in\Theta_{\rm ja}}g\Bigl{(}\text{BP}_{\rm{rx}}(\bm{\theta};\bm{\omega})-\varepsilon_{\rm ja}\Bigr{)}\Biggr{]}\Biggr{\}},$ (13a)
s.t. $\;\|\bm{U}_{q}\tilde{\bm{f}}_{q,r}\|^{2}=1,\;q=1,\ldots,N_{\rm sub}.$ (13b)
@@
Problem (13) is still quite complex, and we propose to find a suboptimal solution by resorting to the block-coordinate ascent method [47]: starting from a feasible point, the objective function is maximized with respect to each of the block of variables, taken in cyclic order, while keeping the other ones fixed at their previous values. The block variables are $\omega_{1}$, $\omega_{2}$,…, $\omega_{D_{\rm ris}}$, and $\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}}$. Therefore, we are faced with $D_{\rm ris}+1$ reduced-complexity sub-problems: $D_{\rm ris}$ sub-problems for the maximization over each RIS phase and one sub-problem for the maximization over the radar beamformers.
@@
As to the RIS phase $\omega_{d}$, a suboptimal solution can be found by resorting to the projected gradient algorithm. In particular, the gradient of the objective function of Problem (13) is reported in (14) at the top of this page, where $g^{\prime}(\chi)=2\max\{0,\chi\}$ is the derivative of $g(\chi)$, and
$\bm{\nabla}_{\!\!\bm{\omega}}$ $=2\sum_{\bm{\theta}\in\Theta_{i}}\sum_{q=1}^{N_{\rm sub}}\Bigl{(}\bm{\omega}^{\mathsf{H}}\bm{B}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}\bm{A}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}+\bm{\omega}^{\mathsf{H}}\bm{A}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}\bm{B}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}\Bigr{)}$
@@
$\quad-2\beta\Bigg{[}\sum_{\bm{\theta}\in\Theta_{i,{\rm sl}}}\!\!g^{\prime}\Bigl{(}\text{BP}(\bm{\theta};\bm{\omega},\{\tilde{\bm{f}}_{q,r}\}_{q=1}^{N_{\rm{sub}}})-\varepsilon_{\rm sl}\Bigr{)}\!\!\sum_{q=1}^{N_{\rm sub}}\!\Bigl{(}\bm{\omega}^{\mathsf{H}}\bm{B}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}\bm{A}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}+\bm{\omega}^{\mathsf{H}}\bm{A}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}\bm{B}_{q}(\bm{\theta};\bm{\omega})\bm{\omega}\Bigr{)}\Bigg{]}$